# citations.bib

### v1

@@ 2.3 Theoretical Guarantees of Shapley Value Approximation
We now provide bounds on the sketching dimension $m$ required to ensure a Shapley value estimator $\widehat{\bm{\phi}}\in\{\bm{\phi}_{\lambda}^{\textup{R}},\bm{\phi}_{\lambda}^{\textup{M}}\}$ satisfies a guarantee
@@
    With replacement: Each of the $m$ rows of $\bm{S}$ are sampled independently. For a given row, a single entry is selected to be nonzero, according the $\mathcal{P}$. The value of this nonzero entry is $1/\sqrt{mp_{S}}$, where $S$ is the index of the nonzero row; see Section A.2.
@@
    Without replacement: For each $S$ we flip a coin that returns heads with probability $q_{S}$. If the coin is heads, we add a row to $\bm{S}$, where the $S$-th entry of the row is nonzero and takes value $1/\sqrt{q_{S}}$. The probabilities $q_{S}$ are chosen based on the $\mathcal{P}$ so that, the dimension of the sketching matrix is equal, on average, to some target value; see Section A.5.
@@
Our main theoretical result, which we prove in Section A.3 using techniques from randomized numerical linear algebra [woodruff2014sketching, Tropp2015, Martinsson2020], is the following:
@@
Define $\bm{P}_{\bm{U}}:=(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})$, and fix $\lambda\in\mathbb{R}$. Sample with or without replacement. Then, the regression estimator satisfies
$m=O\left(\frac{\gamma(\bm{P}_{\bm{U}}\bm{b}_{\lambda})}{\delta\varepsilon^{2}}+\eta\log\left(\frac{d}{\delta}\right)\right)\quad\Longrightarrow\quad\mathbb{P}\left[\left\lVert\bm{\phi}^{*}-\bm{\phi}_{\lambda}^{\textup{R}}\right\rVert<\varepsilon\right]>1-\delta,$
and the matrix multiplication estimator satisfies
$m=O\left(\frac{\gamma(\bm{b}_{\lambda})}{\delta\varepsilon^{2}}\right)\quad\Longrightarrow\quad\mathbb{P}\left[\left\lVert\bm{\phi}^{*}-\bm{\phi}_{\lambda}^{\textup{M}}\right\rVert<\varepsilon\right]>1-\delta.$
A direct computation reveals that $\gamma(\bm{P}_{\bm{U}}\bm{b}_{\lambda})\leq\eta\left\lVert\bm{P}_{\bm{U}}\bm{b}_{\lambda}\right\rVert^{2}\leq\eta\left\lVert\bm{b}_{\lambda}\right\rVert^{2}$, where the first inequality is by the definition of $\eta$ and second inequality is due to the fact that $\bm{P}_{\bm{U}}$ is the orthogonal projector onto the column-span of $\bm{U}$. However, for particular $\bm{b}_{\lambda}$ each of these inequalities may not be sharp.
In Table 1, we provide more refined bounds for the kernel weight, leverage score, and modified row-norm sampling probabilities from Section 2.2. More precise bounds are stated and derived in Section A.4, and we also give bounds for the family of distributions defined in Eq. 2.1 in Remark A.10. Importantly, the bounds for modified row-norm sampling are no worse than leverage scores, but can be up to a factor of $\sqrt{d}$ better in some cases. Furthermore, up to log factors, the bounds for kernel weights are no worse than both leverage scores and modified row-norm sampling, but can be a factor of $d/\log(d)$ or $\sqrt{d}/\log(d)$ better than leverage scores and modified row-norm sampling in some cases, respectively. These observations are formalized in Corollary A.8 and we construct an adversarial model demonstrating better bounds in Appendix E.
@@
Table 1: Bounds (big-$\Theta$) on parameters in Theorem 2.2 for the sampling weights from Section 2.2 (see Corollary A.9). Here $\bm{H}$ is a diagonal matrix for which $\lambda_{\textup{min}}(\bm{H})=\Theta(1/\sqrt{d})$ and $\lambda_{\textup{max}}(\bm{H})=\Theta(1)$. Hence, $\left\lVert\bm{H}\bm{x}\right\rVert/\left\lVert\bm{x}\right\rVert\in[\Theta(1/\sqrt{d}),\Theta(1)]$, so the bounds for kernel sampling are within a $\log(d)$ factor of leverage score sampling in the worst case, but can be better by a factor $d/\log(d)$ in some cases. On the other hand, the bounds for modified $\ell_{2}$ sampling are never worse than leverage score sampling, but can be better by a factor of $\sqrt{d}$ in some cases (see Corollary A.8).
@@ 3 Experiments
Based on our framework, Appendix F describes the pseudo-code of the randomized estimators based on sampling with-replacement Algorithm 1 and without-replacement Algorithm 2. We evaluate these estimators across a range of synthetic and real world settings. Of primary interest is the mean squared error distance from the true Shapley value (normalized: $\texttt{mse}=\lVert\bm{\phi}^{*}-\widehat{\bm{\phi}}\rVert^{2}/\left\lVert\bm{\phi}\right\rVert^{2}$); we explore the convergence of these estimators to the true Shapley Values. We set out to find the best strategy, but our findings reveal that each method has its own merits across different scenarios. A summary of the experiments is provided here, with details deferred to the following sections.
We run experiments on eight popular tabular datasets from the shap library (up to $d=101$) and two image datasets (MNIST $d=64$, and CIFAR-10 $d=3072$), details on each dataset are in Section G.2. In each dataset, we train an XG-Boost model [chen2016xgboost] to compute the exact Shapley values using TreeExplainer class in shap [lundberg2020local2global]. Following [covert2020improving, musco2025provably] we only run our experiments on paired sampling estimators, which are found to be more performant. We report a summary of the experimental findings while leaving detailed experiments to the Appendix G.
In the experiments that follow, [musco2025provably] has been reimplemented to (a) allow the methods to be computed in high dimensions efficiently, (b) to ensure a fair comparison between least squares and matrix-vector approximation by fixing a single $\bm{SZ}$ for both estimators. We run our experiments on an AMD EPYC 7R13 processor with 48 cores per socket, 96 CPUs, and 183GB RAM.
@@ 3.1 Comparisons of Estimators
For each dataset, we choose the first data points of train and test sets, on $80/20$ splits, as baseline and query points for our Shapley estimators respectively. We choose $m=2^{5},2^{6},..,2^{15}$ for larger datasets ($d>12$) and pick specific $m$ values for smaller datasets. We run the experiments on random seeds 0,..,990,..,990 , . . , 99 (numpy and Python’s random) for replicability of results. Exact Shapley values are computed with TreeExplainer on the same baseline; KernelExplainer is run without $\ell_{1}$ regularization. XG-boost regressors and classifiers are fit with $100$ estimators and a maximum tree depth of $10$. We highlight key observations in Fig. 2, where we plot median normalized mean squared errors for $100$, alongside interquantile ranges. Except when specified otherwise, we use $\lambda=\alpha$. Detailed results variations and tables are are in Section H.1. We remark the following findings.
@@
Figure 2: Comparison of performance across different estimators. In (1, top row) estimators use with replacement sampling strategies. In (2,3, central and bottom row) $\bm{SZ}$ is sampled without replacement. In legends, MV refers to matrix-vector and LS to least squares. Dimensions of each datasets are reported with the titles.
@@
  • Matrix-Vector vs. Ordinary Least Squares. We showcase the effectiveness of each Shapley estimator in practice, reporting a comparison between the best performing distribution in Fig. 2 (1, with replacement; 2 without replacement). The clearest separations across methods appears in the comparison between matrix-vector and least squares estimators. We find that least squares based estimators tend to perform better than matrix-vector multiplication. KernelSHAP is generally positioned between these methods in the ranking. This is highlighted in Fig. 2.
@@
  • Comparison Across Distributions. We find that $\ell_{2}$ squared, modified $\ell_{2}$, and the kernel distribution perform comparably. (Fig. 2 (1,2)) $\ell_{2}$ squared outperforming $\ell_{2}$ modified marginally and outperforming kernel more significantly in both least squares and matrix-vector approximations. As discussed in Section 2.2, the choice of sampling distribution affects performance of the estimator; the best choice is problem dependent. In Appendix E, we design a synthetic experiment where $\ell_{2}$ modified distribution significantly outperforms $\ell_{2}$ squared and kernel distributions.
@@
  • With and Without Replacement Sampling Strategies. Sampling strategies perform similarly for matrix-vector. For least-squares, with replacement sampling outperforms without replacement on some datasets. However, this sampling strategy is poorly suited for settings where $m>2^{d}$, since it will perform worse than brute force computation.
@@
  • Comparing $\bm{b}_{0}$ versus $\bm{b}_{\alpha}$ in Matrix-Vector. We compare the performance of the two matrix-vector estimators in Fig. 2 (3). As noted in Appendix B, the unbiased KernelSHAP method of [covert2020improving] uses $\bm{b}_{0}$, while the other methods we have explored use $\bm{b}_{\alpha}$. We find that using $\bm{b}_{\alpha}$ in the estimator leads to better performance.
@@ 3.2 Provably Efficient Methods in High-Dimensions
For high dimensional datasets, we aim to compare the estimators across faithfulness measures, as well as mean squared error. We compute Shapley values on the first 10 data points from the test sets, using the first data point of the training set same baseline, using 80/20 splits. As before, we train a decision tree in order to be able to compare with exact Shapley value computed from the TreeExplainer class. For each method, we compare average normalized mean squared error across test points, computational costs and faithfulness of the explanations. Mean squared error is juxtaposed to time (in seconds) in Fig. 3 1(a,b) to emphasize computational tradeoffs between methods. Faithfulness via both area under the curve (AUC) of insertion and deletion curves in Fig. 3 2(a,b), and Spearman rank correlation between exact and estimated Shapley values (as reported in Fig. 3, 3(a,b)). Detailed experimental results with errors can be found in Section H.2.
Algorithmic Innovations. Approximating Shapley values in high dimensional problems is a challenge. There are two computational bottlenecks in [musco2025provably]: (a) for distributions beyond $\ell_{2}$ squared, combinatorial terms (${d\choose k}$ will be cause overflow/underflow sufficiently large $d$ and middle $k$, (b) even if we are able to compute the term, [musco2025provably] bucket sampling procedure requires binomial sampling from a distribution with support of size (${d\choose k}$. In our Algorithm 2, we overcome both issues for all distributions by (a) avoiding the computation of the combinatorial terms in the probability distributions and weights, and (b) using Poisson approximation of large Binomials to avoid the large support problem. This allows an analysis of our estimators on CIFAR10.
Estimator Performance. In Fig. 3 part 1(a-b), experiments confirm that there is a distinct computational tradeoff between least squares methods and matrix-vector estimators: least squares methods generally requires more time but lead to better approximations for fixed number of samples. Indeed, this discrepancy is accentuated as the number of dimensions increases: possibly due to the time complexity of linear program estimation.
Faithfulness. In Fig. 3 part 2(a-b) and 3(a-b); after 100k samples, we find that for MNIST the ranks of the two estimators are perfectly aligned. For CIFAR-10, there have been significant increases in rank correlation, showcasing the effectiveness of the estimators. In all settings, we find increased fidelity especially as the dimensionality of the problem increases. We note this could be problem dependent. We report AUC curves in Section H.3.
@@ 4 Discussion
Limitations. Computing accurate Shapley values remains a challenge. As with past work, the theoretical bounds we derive for Shapley value estimators depend on quantities involving $\bm{b}_{\lambda}$ (e.g. $\left\lVert\bm{b}_{\lambda}\right\rVert$) which cannot be computed efficiently. As such, they cannot be instantiated by the user. We leave to future work the development of methods that can address this challenge. Moreover, there are several approaches to sampling without replacement and the present work does not provide prescriptions on which to use; this is left to future work.
Future Work. This work promotes trust in the estimation of Shapley values, promoting a responsible use of the estimators in the explainable-AI community. Our theoretical contributions of a unified framework pave the way for development of tailored estimators depending on the observed $\bm{b}_{\lambda}$ values. Future work will design adaptive estimators or develop heuristics tailored to models.

### v2

@@ 2.3 Theoretical Guarantees of Shapley Value Approximation
[…]
@@
    With Replacement: Each of the $m$ rows of $\bm{S}$ are sampled independently. For a given row, a single entry is selected to be nonzero according to $\mathcal{P}$. The value of this nonzero entry is $1/\sqrt{mp_{S}}$, where $S$ is the index of the nonzero row; see Section˜A.2.
@@
    Without Replacement: For each subset $S$, we flip a coin that returns heads with probability $q_{S}$. If the coin is heads, we add a row to $\bm{S}$, where the $S$-th entry of the row is nonzero and takes value $1/\sqrt{q_{S}}$. The probabilities $q_{S}$ are chosen based on the $\mathcal{P}$ so that, the dimension of the sketching matrix is equal, on average, to some target value $m$; see Section˜A.5.
@@
Our main theoretical result, which we prove in Section˜A.3 using techniques from randomized numerical linear algebra [Woo+14, Tro15, MT20], is the following:
@@
Define $\bm{P}_{\bm{U}}:=(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})$, and fix $\lambda\in\mathbb{R}$. Let $m$ denote the sample complexity in the sampling with replacement scenario and the average sample complexity in the sampling without replacement scenario. Then, for the regression estimator,
$m=O\left(\frac{\gamma(\bm{P}_{\bm{U}}\bm{b}_{\lambda})}{\delta\varepsilon^{2}}+\eta\log\left(\frac{d}{\delta}\right)\right)\quad\textnormal{guarantees}\quad\mathbb{P}\left[\left\lVert\bm{\phi}^{*}-\bm{\phi}_{\lambda}^{\textup{R}}\right\rVert<\varepsilon\right]>1-\delta,$
and for the matrix-vector multiplication estimator,
$m=O\left(\frac{\gamma(\bm{b}_{\lambda})}{\delta\varepsilon^{2}}\right)\quad\textnormal{guarantees}\quad\mathbb{P}\left[\left\lVert\bm{\phi}^{*}-\bm{\phi}_{\lambda}^{\textup{M}}\right\rVert<\varepsilon\right]>1-\delta.$
A direct computation reveals that $\gamma(\bm{P}_{\bm{U}}\bm{b}_{\lambda})\leq\eta\left\lVert\bm{P}_{\bm{U}}\bm{b}_{\lambda}\right\rVert^{2}\leq\eta\left\lVert\bm{b}_{\lambda}\right\rVert^{2}$, where the first inequality is by the definition of $\eta$ and second inequality is due to the fact that $\bm{P}_{\bm{U}}$ is the orthogonal projector onto the column-span of $\bm{U}$. However, for a particular $\bm{b}_{\lambda}$, each of these inequalities may not be sharp.
In Table˜1, we provide more refined bounds for the kernel weight, leverage score, and modified row-norm sampling probabilities from Section˜2.2. More precise bounds are stated and derived in Section˜A.4, and we also give bounds for the family of distributions defined in Eq.˜2.1 in Remark˜A.11. Importantly, the bounds for modified row-norm sampling are no worse than leverage scores, but can be up to a factor of $\sqrt{d}$ better in some cases. Furthermore, up to log factors, the bounds for kernel weights are no worse than both leverage scores and modified row-norm sampling, but can be a factor of $d/\log(d)$ or $\sqrt{d}/\log(d)$ better than leverage scores and modified row-norm sampling in some cases, respectively. These observations are formalized in Corollary˜A.9, and we construct an adversarial model demonstrating such an advantage in the sample complexity bounds in Appendix˜E. Intuitively, kernel weights and modified row-norm sampling place a larger importance on subsets of small/large size, as seen from Fig.˜1. As a result, for models where the entries of the vector $\bm{b}_{\lambda}$ or $\bm{P}_{\bm{U}}\bm{b}_{\lambda}$ are concentrated around subsets of small/large size, kernel weights or modified row-norm sampling would perform better than leverage score sampling, which is the key observation we use for constructing the adversarial model in Appendix˜E. It remains to be seen whether kernel weights or modified row-norm sampling scheme provides a sample complexity advantage over leverage scores for models used in practice (such as neural networks), and we leave this as an open question for future research.
@@
Table 1: Bounds (big-$\Theta$) on parameters in Theorem˜2.2 for the sampling weights from Section˜2.2, derived in Corollary˜A.10. $\bm{H}$ is a diagonal matrix defined in Corollary˜A.10 satisfying $\lambda_{\textup{min}}(\bm{H})=\Theta(1/\sqrt{d})$ and $\lambda_{\textup{max}}(\bm{H})=\Theta(1)$, so that $\left\lVert\bm{H}\bm{x}\right\rVert/\left\lVert\bm{x}\right\rVert\in[\Theta(1/\sqrt{d}),\Theta(1)]$. Hence, the bounds for kernel sampling are within a $\log(d)$ factor of leverage score sampling in the worst case, but can be better by a factor $d/\log(d)$ in some cases. On the other hand, the bounds for modified $\ell_{2}$ sampling are never worse than leverage score sampling, but can be better by a factor of $\sqrt{d}$ in some cases (see Corollary˜A.9).
@@ 3 Experiments
Based on our framework, Appendix˜F describes the pseudo-code of the randomized estimators based on sampling with-replacement Algorithm˜1 and without-replacement Algorithm˜2. We evaluate these estimators across a range of synthetic and real world settings. Of primary interest is the mean squared error distance from the true Shapley value (normalized: $\texttt{mse}=\mathbb{E}[\lVert\bm{\phi}^{*}-\widehat{\bm{\phi}}\rVert^{2}]/\left\lVert\bm{\phi}\right\rVert^{2}$); we explore the convergence of these estimators to the true Shapley Values. We set out to find the best strategy, but our findings reveal that each method has its own merits across different scenarios. A summary of the experiments is provided here, with details deferred to the following sections.
In the experiments that follow, [MW25] has been re-implemented to (a) allow the methods to be computed in high dimensions efficiently, and (b) to ensure a fair comparison between regression and matrix-vector multiplication method by fixing a single $\bm{SZ}$ for both estimators. We include results from our implementation of KernelSHAP (regression + kernel weights) as well as the implementation of KernelSHAP from the shap library. This particular implementation includes several additional heuristic optimizations.
We run experiments on eight popular tabular datasets from the shap library (up to $d=101$) and two image datasets (MNIST $d=784$, and CIFAR-10 $d=3072$), details on each dataset are in Section˜G.2. In each dataset, we train an XG-Boost model [CG16] to compute the exact Shapley values using TreeExplainer class in shap [Lun+20]. We report a summary of the experimental findings while leaving detailed experiments to Appendix˜G.
Following [CL20, MW25], we run our experiments using paired sampling, a simple modification of the estimation procedure, which has been observed to improve empirical performance. In paired sampling, when an index $S\subset 2^{[d]}\setminus\{[d],\emptyset\}$ is selected, the compliment $S^{c}=[d]\setminus S$ is also selected. Paired sampling is also used by default in implementation of KernelSHAP from the shap library.
We run our experiments on an AMD EPYC 7R13 processor with 48 cores per socket, 96 CPUs, and 183GB RAM.
@@ 3.1 Comparisons of Estimators
For each dataset, we choose the first data points of the train and the test sets, according to an $80/20$ split, as baseline, and query points for our Shapley estimators respectively. We choose $m=10^{3},10^{4}/2,10^{4},...,10^{6}/2$ for larger datasets ($d>12$) and pick specific values of $m$ for smaller datasets. We run the experiments on random seeds $0,..,99$ (numpy and Python’s random) for replicability of results. Exact Shapley values are computed with TreeExplainer on the same baseline; KernelExplainer is run without $\ell_{1}$ regularization. XG-boost regressors and classifiers are fit with $100$ estimators and a maximum tree depth of $10$. We highlight key observations in Fig.˜2, where we plot median normalized mean squared errors for $100$ random seeds, alongside interquantile ranges. Except when specified otherwise, we use $\lambda=\alpha$. Detailed results, variations and tables are in Section˜H.1. Our findings are as follows.
@@
Figure 2: Comparison of performance across different estimators. In (1, top row) estimators use with replacement sampling strategies. In (2,3, central and bottom row) $\bm{SZ}$ is sampled without replacement. In legends, MV refers to matrix-vector multiplication estimator and LS to regression (least squares) estimator. Dimensions of each datasets are reported with the titles.
@@
  • Matrix-Vector Multiplication vs. Regression Estimator. We showcase the effectiveness of each Shapley value estimator in practice, reporting a comparison between the best performing distribution in Fig.˜2 (1, with replacement; 2 without replacement). The clearest separations across methods appears in the comparison between matrix-vector multiplication and regression estimators. We find that regression estimator tends to perform better than matrix-vector multiplication estimator. KernelSHAP is generally positioned between these methods in the ranking. This is highlighted in Fig.˜2.
@@
  • Comparison Across Distributions. Among the three sampling schemes evaluated, our results in Fig.˜2 (1,2) indicate that the $\ell_{2}$-squared method outperforms modified $\ell_{2}$ marginally, while outperforming kernel distribution more significantly in both regression and matrix-vector multiplication approximations. As discussed in Section˜2.2, the choice of sampling distribution affects the performance of the estimator; with the best choice being problem dependent. In Appendix˜E, we design a synthetic experiment where estimators based on modified $\ell_{2}$ and kernel distributions significantly outperform $\ell_{2}$-squared distribution based estimator.
@@
  • With and Without Replacement Sampling Strategies. Sampling strategies (with and without replacement) perform similarly for the matrix-vector multiplication estimator. For the regression estimator, sampling with replacement outperforms sampling without replacement on some datasets. However, sampling with replacement strategy is poorly suited for settings where $m>2^{d}$, since it will perform worse than brute force computation of Shapley values.
@@
  • Comparing $\lambda$=0 versus $\lambda$$\alpha$ for the Matrix-Vector Multiplication Estimator. We compare the performance of matrix-vector multiplication estimator using $\lambda=0$ and $\lambda=\alpha$ (i.e., $\bm{b}_{0}$ vs $\bm{b}_{\alpha}$) in Fig.˜2 (3). As noted in Appendix˜B, the unbiased KernelSHAP method of [CL20] uses $\lambda=0$, while the other methods we have explored use $\lambda=\alpha$. We find that using $\lambda=\alpha$ in the estimator leads to better performance.
@@ 3.2 Provably Efficient Methods in High-Dimensions
For high dimensional datasets, we aim to compare the estimators across faithfulness measures, as well as mean squared error. We compute Shapley values on the first 10 data points from the test sets, using the first data point of the training set, using 80/20 splits. As before, we train a decision tree in order to be able to compare with exact Shapley value computed from the TreeExplainer class. For each method, we compare average normalized mean squared error across test points, computational costs and faithfulness of the explanations. Mean squared error is juxtaposed with time (in seconds) in Fig.˜3 1(a,b) to emphasize computational tradeoffs between methods. Faithfulness via both area under the curve (AUC) of insertion and deletion curves in Fig.˜3 2(a,b), and Spearman rank correlation between exact and estimated Shapley values (as reported in Fig.˜3, 3(a,b)). Detailed experimental results with errors can be found in Section˜H.2.
Algorithmic Innovations. Approximating Shapley values in high dimensional problems is a challenge. There are two computational bottlenecks in [MW25]: (a) for distributions beyond $\ell_{2}$-squared, combinatorial terms $d\choose k$} will cause overflow/underflow for sufficiently large $d$ and middle $k$ (i.e., $k\sim d/2$), and (b) even if we are able to compute the binomial term, [MW25] bucket sampling procedure requires binomial sampling from a distribution with support of size $d\choose k$}, which can be large. In our Algorithm˜2, we overcome both issues for all distributions by (a) avoiding the computation of the combinatorial terms in the probability distributions and weights, and (b) using Poisson approximation of large binomials to avoid the large support problem. This allows an analysis of our estimators on CIFAR10.
Estimator Performance. In Fig.˜3 part 1(a-b), experiments confirm that regression estimators generally requires less time and lead to better approximations for fixed number of samples compared to matrix-vector multiplication estimator. Indeed, this discrepancy is accentuated as the dimension size increases. The regression estimator produces accurate estimates even when the number of samples is small, improving on all other estimators.
Faithfulness. In Fig.˜3 part 2(a-b) and 3(a-b); after 100k samples, we find that for MNIST, all estimators have similar fidelity, but KernelSHAP has very high rank correlation. This may be due to the fact that KernelSHAP first samples from buckets of size $1$ and $d$, a difference which may be beneficial in this setting. For CIFAR-10, there have been significant increases in rank correlation, showcasing the effectiveness of the estimators. In all settings, we find increased fidelity especially as the dimensionality of the problem increases. We note this could be problem dependent. We report AUC curves in Section˜H.3.
@@ 4 Discussion
Limitations. Computing accurate Shapley values remains a challenge. As with past work, the theoretical bounds we derive for Shapley value estimators depend on quantities involving $\bm{b}_{\lambda}$ (e.g. $\left\lVert\bm{b}_{\lambda}\right\rVert$) which cannot be computed efficiently. As such, they cannot be instantiated by the user. Below, we give a prescription on how this limitation can be mitigated in practice, but leave a thorough study for future research. Also note that there are several approaches to sampling without replacement and the present work does not provide prescriptions on which to use; this is left to future work.
Practical Prescription. As our analysis reveals, the estimators converge in a predictable way with the number of samples ($m$) to the true Shapley value, at the rate of $\sim 1/\sqrt{m}$. Therefore, we can use the estimate from a larger value of $m$ to approximate the error at some $m_{0}\ll m$. As long as $m\gg m_{0}$, the estimate using $m$ samples is a good proxy for the true Shapley values, relative to the error of the estimate using $m_{0}$ samples. We find that this method, while heuristic, gives a good estimate of the error in practice.
Future Work. This work promotes trust in the estimation of Shapley values, promoting a responsible use of the estimators in the explainable-AI community. Our theoretical contributions of a unified framework pave the way for development of tailored estimators depending on the observed entries of $\bm{b}_{\lambda}$, which can be used to adapt the sampling distribution accordingly. Developing such adaptive estimators, as well as their theoretical analysis, is left as an interesting direction for future research.
Acknowledgements
The authors thank Rob Otter and Shaohan Hu for their support and valuable feedback on this project. We also acknowledge our colleagues at the Global Technology Applied Research Center of JPMorganChase, especially Sriram Yechan Gunja and Rajagopal Ganesan, for helpful discussions. We would also like to thank R. Teal Witter and Christopher Musco for providing us the code for LeverageSHAP [MW25].